OpenStudy (anonymous):

State the vertical asymptote of the rational function. f(x)= (x-3)(X+4)/x^2-1

OpenStudy (anonymous):

Factor the denominator. \[x^2 - 1 = (x + 1)(x - 1)\]Now set each factor equal to 0 and solve for x. The reason this works is because there are vertical asymptotes when the denominator is equal to 0. x + 1 = 0 x - 1 = 0 x = ?, ?

OpenStudy (anonymous):

but how does \[x ^{2}-1=(x+1) (x-1)\] ?

OpenStudy (anonymous):

It's the difference in squares pattern. \[(a^2 - b^2) = (a + b)(a - b)\]

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

np :)